# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ Appendix B Appendix: Supplementary Material for Computational Experiments with Adaptive Transport Maps
As indicated in Section 5.2.2 of the main text, in order to assess the robustness of the results obtained from Algorithm 2, “Gathering statistics from transport maps sparsity patterns”, with respect to ordering of the random variables, we carried out the computational experiments described in Section 5.2.2 and, in particular, summarized via Figures 1011 of the main text, using $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ of the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables. Specifically, Algorithm 2 was executed for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ listed in Table 2. The results were then analyzed individually for each permutation, as well as collectively after mapping the results obtained for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$. The outcome of this exercise is summarized next and in Figure 12 of the main text.
@@
Firstly, persistence of patterns of high-valued frequency counts resulting from the computational experiments, as described in Figures 1011 for the permutation $\sigma_{{1}}(\mathcal{{Y}})$ case, was also observed when executing Algorithm 2 for each of the $19$ permutations $\{\sigma_{{m}}(\mathcal{{Y}})\}_{m=2}^{20}$. As one example illustrating this, we display in Figures 1516, for the permutation $\sigma_{{20}}(\mathcal{{Y}})$ case, the analogues of Figures 1011 from the main text. Note, however, that the patterns from Figure 10 versus the corresponding ones from Figure 15 (similarly, Figure 11 versus Figure 16) are different because each were computed for different permutations (namely, $\sigma_{{1}}(\mathcal{{Y}})$ and $\sigma_{{20}}(\mathcal{{Y}})$) of the random variables.
@@
Figure 17: Aggregated sparsity patterns and associated frequency counts from adaptive transport map approximations to data distributions for different subsets of the zero-dose radiation data samples. The assignment of genes to the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 17a17b is that listed as permutation $\sigma_{{1}}(\mathcal{{Y}})$ in Table 2, whereas for Subfigures 17c17d it is that listed as permutation $\sigma_{{20}}(\mathcal{{Y}})$. To facilitate comparison, the frequency counts shown above were mapped to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, as described in Algorithm 3. Specifically, the matrices shown in Subfigures 17b and 17d resulted from line 14 of Algorithm 3. Good qualitative agreement is observed between the high-valued frequency counts for $\mathrm{{P}}_{{0}}^{{(200,\sigma_{{1}})}}$ (Subfigure 17a) and those for $\mathrm{{P}}_{{0}}^{{(200,\sigma_{{20}})}}$ (Subfigure 17c), in spite of quantitative differences (particularly for low-valued frequency counts).
@@
Figure 18: Aggregated sparsity patterns and associated frequency counts from adaptive transport map approximations to data distributions for different subsets of the low-dose radiation data samples. The assignment of genes to the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 18a18b is that listed as permutation $\sigma_{{1}}(\mathcal{{Y}})$ in Table 2, whereas for Subfigures 18c18d it is that listed as permutation $\sigma_{{20}}(\mathcal{{Y}})$. To facilitate comparison, the frequency counts shown above were mapped to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, as described in Algorithm 3. Specifically, the matrices shown in Subfigures 18b and 18d resulted from line 14 of Algorithm 3. Good qualitative agreement is observed between the high-valued frequency counts for $\mathrm{{P}}_{{1}}^{{(200,\sigma_{{1}})}}$ (Subfigure 18a) and those for $\mathrm{{P}}_{{1}}^{{(200,\sigma_{{20}})}}$ (Subfigure 18c), in spite of quantitative differences (particularly for low-valued frequency counts).
Now, to more easily compare and collect the results obtained for the $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, we can map the statistics gathered for each of the permutations to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{j})_{j=1}^{6}$. Precisely, given the matrices $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}$ containing the frequency counts resulting from the transport maps sparsity patterns in a series of $N$ transport maps computed for each class $k\in\{0,1\}$ and, within each class, for the permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ (see line 16 of Algorithm 2), we define new matrices:
  1. 1.
    $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}=\sigma_{{p}}^{-1}\left(\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\right)$: Lower triangular matrix storing the frequency counts from the matrix $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\,$, rearranged after mapping from the ordering of the random variables corresponding to the permutation $\sigma_{{p}}(\mathcal{{Y}})$ to that for the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{j})_{j=1}^{6}$, that is $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)=\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\left(\min(\sigma_{{p}}(j),\sigma_{{p}}(i)),\max(\sigma_{{p}}(j),\sigma_{{p}}(i)\right)$ (see lines 1214 of Algorithm 3). An entry $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)$ in the lower triangular portion of the matrix $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ thus contains a frequency count reflecting a dependency “strength” between the random variables $\mathcal{{Y}}_{j}$ and $\mathcal{{Y}}_{i}$, obtained after computing a series of $N$ transport maps with an ordering $\sigma_{{p}}(\mathcal{{Y}})$ of the random variables, after being mapped back (i.e., rearranged) so that it can be compared with the matrix of frequency counts obtained from the computation of $N$ transport maps with an ordering $\sigma_{{1}}(\mathcal{{Y}})$ of the random variables or another matrix of frequency counts corresponding to another random variable ordering $\sigma_{{q}}(\mathcal{{Y}})$ after an analogous “normalization” (via the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$). Such “normalization” simplifies comparison of results between separate sets of runs for each one of the $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$. An example is provided in Figures 1718 for the instance $N\!=\!200$ of runs for permutations $\sigma_{{1}}(\mathcal{{Y}})$ and $\sigma_{{20}}(\mathcal{{Y}})$, where one can observe (qualitatively) good agreement between the resulting patterns of high-valued frequency counts. We observed (qualitatively) the same patterns with each one of the remaining permutations of random variables in Table 2, after computing a series of $N\!=\!200$ transport maps and subsequent “normalization” (in the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ – see line 14 of Algorithm 3). This shows robustness of the numerical method and its invariance with respect to the action of elements of the group of permutations $\mathcal{S}_{6}$ when identifying the dependence structure of the set of random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ as inferred from the sparsity patterns of the computed transport maps or, equivalently, from the structure of the corresponding estimated target densities.
  2. 2.
    $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$: Lower triangular matrix accumulating the frequency counts from the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ defined above. That is, for a given value of $N$ runs and for each class $k\in\{0,1\}$,
    where $n_{\mathrm{perm}}\!\!=\!\!20$ denotes the number of permutations and $\widetilde{N}=N\times n_{\mathrm{perm}}$. This summarizes frequency counts resulting from runs for all permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, collectively, as illustrated in Figure 12 of the main text.
The above described matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ and $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$ are created as outlined in Algorithm 3.
1:Let $m(=6)$ denote the number of random variables
@@
2:Let $n_{\mathrm{perm}}(=20)$ denote the number of permutations $\sigma_{{p}}(\mathcal{{Y}})$ of the random variables
3:Let $N$ denote a given number of runs
4:Let $\widetilde{N}=N\times n_{\mathrm{perm}}$
5:Let $K(=2)$ denote the number of classes
6:for $k=0$ to $K-1$ do
7:     Initialize $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$ to the $m\times m$ zero matrix
8:     for $p=1$ to $n_{\mathrm{perm}}$ do
9:         Initialize $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ to the $m\times m$ zero matrix
10:         for $j=1$ to $m$ do
11:              for $i=1$ to $j$ do
@@
14:                  $ȷ~ı\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)\ =\ \mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}(\tilde{\mbox{{\j}}},\tilde{\mbox{{\i}}})$triangleright frequency count reflecting dependency
15:                                                                   “strength” between random variables
16:                                                                   $\mathcal{{Y}}_{j}$ and $\mathcal{{Y}}_{i}$, of $N$ runs for permutation
17:                                                                   $\sigma_{{p}}(\mathcal{{Y}})$
18:                  $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}(j,i)\ =\ \mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}(j,i)+\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)$triangleright accumulate frequency counts
19:                                                                                from line 14
20:              end for
21:         end for
22:     end for
23:end for
Algorithm 3

### v4

@@ Appendix B Appendix: Supplementary Material for Computational Experiments with Adaptive Transport Maps
As indicated in Section 5.2.2 of the main text, in order to assess the robustness of the results obtained from Algorithm 2 with respect to ordering of the random variables, we carried out the computational experiments described in Section 5.2.2 and, in particular, summarized via Figures 1011 of the main text, using $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ of the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables. Specifically, Algorithm 2 was executed for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ listed in Table 2. The results were then analyzed individually for each permutation, as well as collectively after mapping the results obtained for each of the $20$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$. The outcome of this exercise is summarized next and in Figure 12 of the main text.
Firstly, persistence of patterns of high-valued frequency counts resulting from the computational experiments, as described in Figures 1011 for the permutation $\sigma_{{1}}(\mathcal{{Y}})$ case, was also observed when executing Algorithm 2 for each of the $19$ permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{m=2}^{20}$. As one example illustrating this, we display in Figures 1516, for the permutation $\sigma_{{20}}(\mathcal{{Y}})$ case, the analogues of Figures 1011 from the main text. Note, however, that the patterns from Figure 10 versus the corresponding ones from Figure 15 (similarly, Figure 11 versus Figure 16) are different because each were computed for different permutations (namely, $\sigma_{{1}}(\mathcal{{Y}})$ and $\sigma_{{20}}(\mathcal{{Y}})$) of the random variables.
So, to more easily compare and collect the results obtained for the $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, we map the statistics gathered for each of the permutations to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{j})_{j=1}^{6}$. Precisely, given the matrices $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}$ (see line 16 of Algorithm 2) containing the frequency counts resulting from the transport maps sparsity patterns in a series of $N$ adaptive transport maps learned for each class $k\in\{0,1\}$ and, within each class, for the permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$, we define new matrices, as described next.
First, for each of the matrices $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}$ (from line 16 of Algorithm 2), we create a matrix $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}=\sigma_{{p}}^{-1}\left(\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\right)$,which is a lower triangular matrix storing the frequency counts from the given matrix $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\,$, rearranged after mapping from the ordering of the random variables corresponding to the permutation $\sigma_{{p}}(\mathcal{{Y}})$ to that for the identity permutation $\sigma_{{1}}(\mathcal{{Y}})=(\mathcal{{Y}}_{j})_{j=1}^{6}$. That is,
$\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)$ $=$ $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}\left(\min(\sigma_{{p}}(j),\sigma_{{p}}(i)),\max(\sigma_{{p}}(j),\sigma_{{p}}(i)\right)$ (49)
(see lines 1719 of Algorithm 3). An entry $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)$ in the lower triangular portion of the matrix $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ thus contains a frequency count reflecting a dependency “strength” between the random variables $\mathcal{{Y}}_{j}$ and $\mathcal{{Y}}_{i}$, obtained after learning a series of $N$ transport maps with the (particular) ordering $\sigma_{{p}}(\mathcal{{Y}})$ of the random variables, after being mapped back (i.e., rearranged) so that it can be compared with the matrix of frequency counts obtained from the learning of $N$ transport maps with the ordering $\sigma_{{1}}(\mathcal{{Y}})$ of the random variables, or with another matrix of frequency counts corresponding to another ordering $\sigma_{{q}}(\mathcal{{Y}})$, $q\neq p,$ of the random variables, after an analogous rearrangement (via the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$) from the ordering $\sigma_{{q}}(\mathcal{{Y}})$ to the ordering $\sigma_{{1}}(\mathcal{{Y}})$ of the random variables. This rearrangement simplifies comparison of results between separate sets of runs for each one of the $20$ different permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$.
An example of such rearrangement is provided in Figures 1718 for the instance $N\!=\!200$ of runs for permutations $\sigma_{{1}}(\mathcal{{Y}})$ and $\sigma_{{20}}(\mathcal{{Y}})$, where one can observe (qualitatively) good agreement between the resulting patterns of high-valued frequency counts. We observed (qualitatively) the same patterns with each one of the remaining permutations of random variables listed in Table 2, after learning series of $N\!\in\!\{10,20,50,100,200\}$ adaptive transport maps and the subsequent rearrangement (in the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ – see line 19 of Algorithm 3) to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$. This shows robustness of the numerical method and its invariance with respect to the action of elements of the group of permutations $\mathcal{S}_{6}$ when identifying the dependence structure of the set of random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ as inferred from the sparsity patterns of the learned adaptive transport maps or, equivalently, from the structure of the corresponding estimated target densities.
Finally, for each class $k\in\{0,1\}$, we create a matrix $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$, which is a lower triangular matrix accumulating the frequency counts from the matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ defined above. That is, for a given value $N$ of the number of transport maps in a series of learned adaptive transport maps and for each class $k\in\{0,1\}$,
where $n_{\mathrm{perm}}$ denotes the number of permutations on the ordering of the random variables (see Table 2) and $\widetilde{N}=N\times n_{\mathrm{perm}}$. This summarizes frequency counts resulting from execution of Algorithm 2 for all permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{20}$ listed in Table 2, collectively (see line 23 from Algorithm 3), as illustrated in Figure 12 of the main text. The above described matrices $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ (49) and $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$ (50) are created as outlined in Algorithm 3.
@@
Figure 17: Aggregated sparsity patterns and associated frequency counts from adaptive transport map approximations to data distributions for different subsets of the zero-dose radiation data samples. The assignment of genes to the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 17a17b is that listed as permutation $\sigma_{{1}}(\mathcal{{Y}})$ in Table 2, whereas for Subfigures 17c17d it is that listed as permutation $\sigma_{{20}}(\mathcal{{Y}})$. To facilitate comparison, the frequency counts shown above were mapped to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, as described in Algorithm 3. Specifically, the matrices shown in Subfigures 17b and 17d resulted from line 19 of Algorithm 3. Good qualitative agreement is observed between the high-valued frequency counts for $\mathrm{{P}}_{{0}}^{{(200,\sigma_{{1}})}}$ (Subfigure 17a) and those for $\mathrm{{P}}_{{0}}^{{(200,\sigma_{{20}})}}$ (Subfigure 17c), in spite of quantitative differences (particularly for low-valued frequency counts).
@@
Figure 18: Aggregated sparsity patterns and associated frequency counts from adaptive transport map approximations to data distributions for different subsets of the low-dose radiation data samples. The assignment of genes to the ordered set $\mathcal{{Y}}\equiv(\mathcal{{Y}}_{j})_{j=1}^{6}$ of random variables for the computational experiments summarized in Subfigures 18a18b is that listed as permutation $\sigma_{{1}}(\mathcal{{Y}})$ in Table 2, whereas for Subfigures 18c18d it is that listed as permutation $\sigma_{{20}}(\mathcal{{Y}})$. To facilitate comparison, the frequency counts shown above were mapped to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$, as described in Algorithm 3. Specifically, the matrices shown in Subfigures 18b and 18d resulted from line 19 of Algorithm 3. Good qualitative agreement is observed between the high-valued frequency counts for $\mathrm{{P}}_{{1}}^{{(200,\sigma_{{1}})}}$ (Subfigure 18a) and those for $\mathrm{{P}}_{{1}}^{{(200,\sigma_{{20}})}}$ (Subfigure 18c), in spite of quantitative differences (particularly for low-valued frequency counts).
1:Notation:
2:    $m\geq 1$: number of random variables
@@
3:    $n_{\mathrm{perm}}$: number of permutations $\sigma_{{p}}(\mathcal{{Y}})$ of the random variables
4:    $N$: number of transport maps in a series of learned adaptive transport maps
5:    $\widetilde{N}$: the value $N\times n_{\mathrm{perm}}$
6:    $K$: number of classes
7:    $\mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}$: $m\times m$ lower triangular matrix storing frequency counts from sparsity patterns for
8:        $N$ adaptive transport maps learned for class $k$ and permutation $\sigma_{{p}}(\mathcal{{Y}})$
9:        (see line 16 of Algorithm 2)
10:Algorithm:
11:for $k=0$ to $K-1$ do
12:     Initialize $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}$ to the $m\times m$ zero matrix
13:     for $p=1$ to $n_{\mathrm{perm}}$ do
14:         Initialize $\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}$ to the $m\times m$ zero matrix
15:         for $j=1$ to $m$ do
16:              for $i=1$ to $j$ do
@@
19:                  $ȷ~ı\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)\ =\ \mathrm{{B}}_{{k}}^{{(N,\sigma_{{p}})}}(\tilde{\mbox{{\j}}},\tilde{\mbox{{\i}}})$triangleright frequency count reflecting dependency
20:                                                                                              “strength” between random variables
21:                                                                                              $\mathcal{{Y}}_{j}$ and $\mathcal{{Y}}_{i}$, out of $N$ runs for
22:                                                                                              permutation $\sigma_{{p}}(\mathcal{{Y}})$
23:                  $\mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}(j,i)\ =\ \mathrm{{Q}}_{{k}}^{{(\widetilde{N})}}(j,i)+\mathrm{{P}}_{{k}}^{{(N,\sigma_{{p}})}}(j,i)$triangleright accumulate frequency counts
24:                                                                                                             from line 19
25:              end for
26:         end for
27:     end for
28:end for
Algorithm 3 Mapping statistics gathered from results for a number of permutations $\{\sigma_{{p}}(\mathcal{{Y}})\}_{p=1}^{n_{\mathrm{perm}}}$ of the random variables to the random variable ordering corresponding to the identity permutation $\sigma_{{1}}(\mathcal{{Y}})$